Another lockdown challenge to recharge your mind.
A teacher asked a smart girl question about her age!
Teacher how old are u?
Girl: I don't know but i know my mother's age twice my age.
Teacher: Then how old is your mother?
Girl: I don't know but i know my father is 5 years older than my mother.
Teacher: How old is your father?
Girl: I don't know but i know the sum total of all our ages are 100.
Question:
1. How old is the girl
2. How old is the mum
3. How old is the dad Please Show your working. 😎🤓😎😀

Answer by mananth(16946)


Another lockdown challenge to recharge your mind.
A teacher asked a smart girl question about her age!
Teacher how old are u?
my age x
Girl: I don't know but i know my mother's age twice my age.
Mother's age =2x
Teacher: Then how old is your mother?
Girl: I don't know but i know my father is 5 years older than my mother.
Teacher: How old is your father?
Father's age = 2x+5
Girl: I don't know but i know the sum total of all our ages are 100.
x+2x+2x+5 =100
5x=95
x=19 ---girl's age
Question:
1. How old is the girl
19 years
2. How old is the mum
38 years
3. How old is the dad
38+5= 43 years
43+38+19 =100